# Simplicity in Complexity

### v1

@@ Introduction
To choose hand-crafted features, we start from the observation that features that fragment images in meaningful ways tend to estimate complexity relatively well (for example, clutter in (?, ?) or the number of regions in (?, ?)). We therefore leverage the capabilities of state of the art (SOTA) image segmentation models to extract relevant segments from the image at multiple spatial granularities. Such models are the closest existing approximations to how humans represent scenes for two main reasons: first, the models are trained using a vast amount of annotations from several humans and hence reflect relevant inductive biases, and second, the architecture of CNNs and transformers are inspired by the human visual processing systems and generalize surprisingly well to unseen images. With the help of such models, we obtain semantically consistent segments at different spatial granularities relevant to perceptual image processing (?, ?). We then derive from them the core components of perceived complexity.
@@
Finding Segments using a Foundational Segmentation Model
@@
We found the nameable class instances in an image using FC-CLIP (?, ?). FC-CLIP is an open-vocabulary panoptic segmentation algorithm that can find multiple instances of each class, and achieves SOTA performance (?, ?). We use panoptic semantic segmentation to predict classes because multi-scale methods like Semantic SAM (?, ?) produced many false positives. We set all parameters of FC-CLIP to default and evaluated the number of detected classes per image (num_class).
@@
We estimate subjective complexity using multiple linear regression. A preliminary examination showed that subjective complexity scales roughly linearly with $\sqrt{num\_seg}$ and $\sqrt{num\_class}$, hence, we apply a square-root transformation to our features. We used the statsmodels OLS function in Python to fit our models. We perform 3-fold cross-validation $M$ times, where $M$ is larger for smaller datasets, and report the average Spearman correlation over all test sets. We compare our models to six baselines from previous work. These baselines include three handcrafted feature-based baselines—two from (?, ?)–Corchs 1, comprising of their 3 best features M8, M5 and M10 (only tested on RSIVL since we were unable to implement M8 (number of regions) to apply it for other datasets), Corchs 2 comprising of 10 features M1 to M11 (excluding M8) and one baseline from (?, ?) comprising of their clutter and patch-wise symmetry measures. The other three are CNN baselines, namely the supervised method from (?, ?), the transfer-learning method from (?, ?) and the supervised method from (?, ?).
@@ Results
It performs similarly to the supervised neural network from (?, ?). The exceptions are the test datasets from the same paper and Savoias Art. This neural network directly learns a high-dimensional mapping from image to complexity, thereby discovering features that best predict complexity in a supervised way. We show that in many cases, this high-dimensional relationship can be distilled down to simply the number of segments and named instances in the image. Hence, we provide evidence that complexity is computable from segmentation features, rather than requiring features that are explicitly optimized for complexity.
@@
Figure 3 shows the mean and standard deviation of the ground truth subjective complexity for images for each bin of $\sqrt{\textit{num\_seg}}$ and $\sqrt{\textit{num\_class}}$ on an example image-set, IC9600 Scenes. In general, and as expected, mean ground truth complexity increases with increasing $\sqrt{\textit{num\_class}}$ and increasing $\sqrt{\textit{num\_seg}}$ (also well-matched to predictions) showing that a complex image is one with both a large number of segments and classes. The standard deviation of the subjective complexity, which contributes markedly to the prediction error, is particularly high in the bins with the greatest and least $\sqrt{\textit{num\_class}}$. This suggests that FC-CLIP might over- or under-predict classes. The largest discrepancies lie in the bin with the highest $\sqrt{\textit{num\_seg}}$ and the lowest $\sqrt{\textit{num\_class}}$. Here, FC-CLIP often fails to find any nameable segments at all.
@@
Refer to caption (a) G: 16, P: 56, Symm: 97 Refer to caption (b) num_seg: 204 (78.5) Refer to caption (c) num_class: 37 (87) Refer to caption (d) G: 56, P: 83, Symm: 53 Refer to caption (e) num_seg: 538 (100) Refer to caption (f) num_class: 3 (9.3)
@@
The statistics of segments and classes at multiple granularities explain most of the variance in the datasets we tested. However, the structure in the image, i.e., the spatial and functional relationships between elements is also known to be an important contributor to complexity ((?, ?, ?), Gestalt theory of perception). Indeed, we find that segment statistics alone are not enough to adequately explain complexity judgments in two other image-sets: VISC and Savoias Interior Design. Figure 5 shows an example from each dataset with the highest prediction errors. In each case, our model over-predicts complexity because SAM finds too many segments without accounting for the fact that many segments are arranged in a spatial pattern (books in the top row and windows in the bottom row). Further, num_class does not contribute to reducing complexity in such cases, either because it is also high (since the uniqueness of classes is not accounted for), or because the weight of the num_class term is learned to be low (for example in VISC) We see that both of these images have high patch-symmetry (defined in (?, ?)). Figure 5 illustrates a significant, positive correlation between patch-symmetry and model error (prediction minus ground truth), showing that our model tends to over-predict when the image is more spatially symmetric, i.e. has more spatial structure. Table 2 shows that when patch_symmetry is added as a feature to the regression, our model improves in Spearman correlation by atleast 0.12, and becomes competitive with most baselines.
@@ Discussion
Our model performance was generally comparable to the supervised neural network of (?, ?), which was trained on a large dataset to directly predict complexity. The difference in performance can be explained by the fact that the neural network potentially utilizes many more than two features and conditionally chooses those features based on the context and distribution of images. However, we have shown that only two features can explain complexity equally well on multiple datasets and domains, elucidating a simpler view of complexity.
@@
We modeled subjective complexity ratings which was the mean rating across multiple raters. However, complexity judgments are known to vary across both individuals or groups (age, cultures, etc.) (?, ?). For the art datasets, the complexity ratings were given by art-novices and would likely differ significantly from ratings of art-experts (?, ?, ?). These individual differences could be caused by differences in the segments people perceive (the regions of an image they consider to be part of the same segment). For example, different individuals may segment at different granularities. Individual differences can also be caused by the mapping from the perceived segments to complexity. Explicitly accounting for individual variability using subject-specific data (for example by fine-tuning the regression or segmentation models) will be an important part of future work.
In conclusion, we develop a parsimonious and interpretable account of human perceived complexity in naturalistic images using segmentation-based methods, showing that complexity can be surprisingly simple given the right image representations.

### v2

@@ Introduction
To choose hand-crafted features, we start from the observation that features that fragment images in meaningful ways tend to estimate complexity relatively well (for example, clutter in (?, ?) or the number of regions in (?, ?)). We therefore leverage the capabilities of state of the art (SOTA) image segmentation models to extract relevant segments from the image at multiple spatial granularities. Such models are the closest existing approximations to how humans represent scenes for two main reasons: first, the models are trained using a vast amount of annotations from several humans and hence reflect relevant inductive biases, and second, the architecture of CNNs and transformers are loosely inspired by the human visual processing systems and generalize surprisingly well to unseen images. With the help of such models, we obtain semantically consistent segments at different spatial granularities relevant to perceptual image processing (?, ?). We then derive from them the core components of perceived complexity.
@@
Finding Segments using a Foundation Segmentation Model
@@
We found the nameable class instances in an image using FC-CLIP (?, ?). FC-CLIP is an open-vocabulary panoptic segmentation algorithm that can find multiple instances of each class, and achieves SOTA performance (?, ?). We use panoptic semantic segmentation to predict classes because multi-scale methods like Semantic SAM (?, ?) produced many false positives. We set all parameters of FC-CLIP to default and evaluated the number of detected classes (including repeated classes) per image (num_class).
@@
We estimate subjective complexity using multiple linear regression. A preliminary examination showed that subjective complexity scales roughly linearly with $\sqrt{num\_seg}$ and $\sqrt{num\_class}$, hence, we apply a square-root transformation to our features. We used the statsmodels OLS function in Python to fit multiple linear regression on each image-set. We perform 3-fold cross-validation $M$ times, where $M$ is larger for smaller image-sets, and report the average Spearman correlation over all test sets. We compare our models to six baselines from previous work. These baselines include three handcrafted feature-based baselines—two from (?, ?)–Corchs 1, comprising of their 3 best features M8, M5 and M10 (only tested on RSIVL since we were unable to implement M8 (number of regions) to apply it for other datasets), Corchs 2 comprising of 10 features M1 to M11 (excluding M8) and one baseline from (?, ?) comprising of their clutter and patch-wise symmetry measures. The other three are CNN baselines, namely the supervised method from (?, ?), the transfer-learning method from (?, ?) and the supervised method from (?, ?).
@@ Results
Our model performs similarly to the supervised neural network from (?, ?). The exceptions are the test datasets from the same paper and Savoias Art. The neural network from (?, ?) directly learns a high-dimensional mapping from image to complexity, thereby discovering features that best predict complexity in a supervised way. We show that in many cases, this high-dimensional relationship can be distilled down to simply the number of segments and named instances in the image. For instance, we find high correlation between the predictions of our model and that of Feng et al. (?) (for example, $r=0.93$ on RSIVL and $r=0.85$ on IC9600 scenes). Moreover, combining the regressors of our model and the Feng et al. model (?) did not increase performance above the best model significantly (with the Savoias Scenes dataset being a notable exception achieving a correlation of $r=0.85$). Hence, we provide evidence that complexity is computable from segmentation features, rather than requiring features that are explicitly optimized for complexity.
@@
Figure 3 shows the mean and standard deviation of the ground truth subjective complexity for images for each bin of $\sqrt{\textit{num\_seg}}$ and $\sqrt{\textit{num\_class}}$ on an example image-set, IC9600 Scenes. In general, and as expected, mean ground truth complexity increases with increasing $\sqrt{\textit{num\_class}}$ and increasing $\sqrt{\textit{num\_seg}}$ (also well-matched to predictions) showing that a complex image is one with both a large number of segments and classes. The standard deviation of subjective complexity, which contributes markedly to the prediction error, is particularly high in the bins with the greatest and least $\sqrt{\textit{num\_class}}$. This suggests that FC-CLIP might over- or under-predict classes. The largest discrepancies lie in the bin with the highest $\sqrt{\textit{num\_seg}}$ and the lowest $\sqrt{\textit{num\_class}}$. Here, FC-CLIP often fails to find any nameable segments at all.
@@
(a) G: 16, P: 56, Symm: 97
(b) num_seg: 204 (78.5)
(c) num_class: 37 (87)
(d) G: 56, P: 83, Symm: 53
(e) num_seg: 538 (100)
(f) num_class: 3 (9.3)
@@
The statistics of segments and classes at multiple granularities explain most of the variance in the datasets we tested. However, the structure in the image, i.e., the spatial and functional relationships between elements is also known to be an important contributor to complexity ((?, ?, ?), Gestalt theory of perception). Indeed, we find that segment statistics alone are not enough to adequately explain complexity judgments in two other image-sets: VISC and Savoias Interior Design. Figure 5 shows an example from each dataset with the highest prediction errors. In each case, our model over-predicts complexity because SAM finds too many segments without accounting for the fact that many segments are arranged in a spatial pattern (books in the top row and windows in the bottom row). Further, num_class does not contribute to reducing complexity in such cases, either because it is also high (since the uniqueness of classes is not accounted for), or because the weight of the num_class term is learned to be low (for example in VISC) We see that both of these images have high patch-symmetry, a measure of spatial regularity based on the average reflection symmetry of local patches of different sizes in the image (see (?, ?) for details). Figure 5 illustrates a significant, positive correlation between patch-symmetry and model error (prediction minus ground truth), showing that our model tends to over-predict when the image is more spatially symmetric, i.e. has more spatial structure. Table 2 shows that when patch_symmetry is added as a feature to the regression, our model improves in Spearman correlation by atleast 0.12, and becomes competitive with most baselines.
@@ Discussion
Our model performance was generally comparable to the supervised neural network of (?, ?), which was trained on a large dataset to directly predict complexity. The difference in performance can be attributed to the neural network potentially utilizing many more than two features and conditionally choosing them based on the context and distribution of images. However, we show that only two features can explain complexity equally well on multiple datasets and domains, elucidating a simpler view of complexity.
@@
We modeled subjective complexity ratings which represent the mean rating across multiple raters. However, complexity judgments are known to vary across both individuals or groups (age, cultures, etc.) (?, ?). For the art datasets, the complexity ratings were given by art-novices and would likely differ significantly from ratings of art-experts (?, ?, ?). These individual differences could be caused by differences in the segments people perceive (the regions of an image they consider to be part of the same segment). For example, different individuals may segment at different granularities. Individual differences can also be caused by the mapping from the perceived segments to complexity. Explicitly accounting for individual variability using subject-specific data (for example by fine-tuning the regression or segmentation models) will be an important part of future work.
In conclusion, we develop a parsimonious and interpretable account of subjective complexity in naturalistic images using segmentation-based methods, showing that complexity can be surprisingly simple given the right image representations.